OpenStudy (aaronandyson):

Why cant we divide by 0?

OpenStudy (mayankdevnani):

0 times anything is 0

OpenStudy (aaronandyson):

I know that, I'm asking that why is "n" divided by 0 equal to infinity/not defined.

OpenStudy (zzr0ck3r):

Suppose you could. \(\dfrac{a}{0}=k\) where \(k\) is some number and \(a\ne 0\). Multiply both sides by \(0\). \(a=0\) YIKES!

OpenStudy (zzr0ck3r):

hehe

OpenStudy (zzr0ck3r):

Look at these numbers \(\dfrac{1}{1},\dfrac{1}{0.01},\dfrac{1}{0.001},\dfrac{1}{0.00001},\dfrac{1}{0.00000000001},\dfrac{1}{0.00000000000000000001}\)

ganeshie8 (ganeshie8):

It is definitely not a law passed by your government to prohibit division by 0. It helps to notice that 1/x approaches positive infinity as x approaches 0 from the right side of the origin, as shown by zzr above.